OpenStudy (anonymous):

what is the ratio of 4m to 8m?

OpenStudy (anonymous):

4/8=2

OpenStudy (anonymous):

4m:8m =1m:2m

OpenStudy (anonymous):

1 to 2

OpenStudy (anonymous):

4m to 8m = 4m / 8m = 4/8 = 1/2 = 1:2

OpenStudy (anonymous):

1/2=0.5

OpenStudy (anonymous):

@Farmer_Wu It is expected to be in ratio form. So it would be 0.5:1, but 1:2 is good too

OpenStudy (anonymous):

yes,it will have many way to wr

OpenStudy (anonymous):

2